OpenStudy (anonymous):

solve the square root of x+4 = the square root of x plus the square root of 2

OpenStudy (anonymous):

\[\sqrt{x+4}=\sqrt{x}+\sqrt2\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

answer not obvious for this one square both sides first the left will be \(x+4\) the right you have to multiply out \((\sqrt{x}+\sqrt2)(\sqrt{x}+\sqrt2)\)